OpenStudy (anonymous):

( x- 7)^2 = 45 how would we solve this for x ?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

We need to use formula of expansion ( a - b )^ = a^2-2ab+b^2 can you try ?

OpenStudy (nikato):

But you can simplify sqrt45 further

OpenStudy (anonymous):

can x be a negative?

hartnn (hartnn):

rather try taking square root on both sides as your first step

hartnn (hartnn):

and there are 2 values of x , not just 1

hartnn (hartnn):

\(\large (x-7) = \pm \sqrt{45} \\ ...\)

OpenStudy (jadeishere):

x2−14x+49=45 so it's a quadratic equation meaning.. quadratic formula >.<

hartnn (hartnn):

all that isn't required! just add 7 on both sides of the equation i posted..

OpenStudy (jadeishere):

Are you sure? O.O

hartnn (hartnn):

yep :)

OpenStudy (jadeishere):

I already know the answers, so if I get the same answer with that that'd be really easy. Just add 7 to both sides?

hartnn (hartnn):

absolutely, add 7 to both sides of \(\large (x-7) = \pm \sqrt{45} \\ \)